OpenStudy (anonymous):

Express in trams of sums and differences of logarithms loga^5x5yz^6 help :(

OpenStudy (anonymous):

\[\log _{a}^{5x5}yz ^{6}\]

OpenStudy (solomonzelman):

\[\log(a^b)=(b)\log(a)\]\[\log(a \times b)=log(a)+log(b)\] I don't get this, can you draw it?

OpenStudy (anonymous):

\[\log _{a}^{5x5}yz ^{6}\]

OpenStudy (solomonzelman):

\[\log_a~^{5x5}yz^6\]

OpenStudy (solomonzelman):

\[5x5?~~maybe~~5x^5?\]

OpenStudy (anonymous):

your right

OpenStudy (anonymous):

\[loga ^{5x ^{5}}yz ^{6}\]

OpenStudy (solomonzelman):

\[\huge\color{blue}{\log(a)^{5x^5}yz^6~~~~~this?}\]

OpenStudy (anonymous):

right

OpenStudy (solomonzelman):

\[5Log(5x) \times Log(a)+6Log(yz)~~~~~~explanations?\]

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

I used \[Log(a \times b)=\log(a)+\log(b)\]to get\[\huge\color{blue}{\huge {\bbox[5pt,cyan,border:2px solid purple]{Log(a)^{5x^5} \times yz^6=Log(a)^{5x^5}+Log(yz^6)}}}\]

OpenStudy (solomonzelman):

then\[loga^b=a \log b\]

OpenStudy (solomonzelman):

\[\huge\color{blue}{\huge {\bbox[5pt,cyan,border:2px solid purple]{5x^5Log(a)+6Log(yz)}}}\]

OpenStudy (solomonzelman):

then the answer

OpenStudy (anonymous):

25x log a + 6 log yz

OpenStudy (solomonzelman):

was the exponent 5 going on 5x entirely, 5 as well or just x?

OpenStudy (anonymous):

just x